OpenStudy (anonymous):

Marcus jogs 4 mi around a track at an average rate of r mi/h. Then he cools down by walking 1 mi at a rate 3mi/h slower. His whole workout lasts 1 h. At what rates did Marcus jog and walk?

OpenStudy (anonymous):

this is a time total motion problem--- p/r=4 1st p  sorry if you can't understand the second drawing so i am including it in a different form below 4/r+1/r-3=1hour (r) (r-3) taken from each (4/r) and so the two r's cancel + (1/r-3) 2 r-3's cancel solve equasion 4r-8+r=r^2-3r should get the rates for you hope this helps